Find (q∘r)(x).
q(x)=x+2
r(x)=x2

Answer by stanbon(75887)

Find (q∘r)(x).
q(x)=x+2
r(x) = x^2
--------------------
qor(x) = q[r(x)] = q[x^2] = x^2+2
-----------------
Cheers,
Stan H.